OpenStudy (anonymous):

2x + 6 ≥ –4 A. x ≥ –5 B. x ≥ –8 C. x ≤ –5 D. x ≤ –8

OpenStudy (dmezzullo):

-6 on both sides

OpenStudy (shaik0124):

2x+6-6>=-4-6 2x>=-10 divide by 2 on both sides now x>=-5

OpenStudy (shaik0124):

@madi understood

OpenStudy (shaik0124):

A is option

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

3x – 6 = –2x + 9 A. 3/5 B. 5/2 C. 3 D. 15